Find the equation of the normal line to the hyperbola x^2 − 2y^2 = 1 at the point (− 3,2) . (Use implicit differentiation).
Answer #1

o a hyperbola at a potrd perpendicular to Through the point the tangent a at 1 and the normal to Equation of line is the equaمام 3. U At (-32) islope of normal = -2(2) (-3) equation of normal = y-Yam (X-*) (4 - 2) = 쓸 4 (x+3) 3 =h 4x 3 + 6